[Brahms-l] Shift report 20050517 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 17 2005 - 00:02:09 EDT
Shifter leader was rw
Shifters were 


1. Beam 5hour, from 16:30-19;00, 21:30-the of the shift

2. Data run 14698-14705

3. All above runs have HV problem. 
   Mainframe BB3 had some hardware problem (according to Fleming) 
   and cannot reset the proper HVs for those connected to that 
   mainframe. Those are BBR31-35 BBL37-44, CCL1-19. 

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 17 00:02:26 2005

This archive was generated by hypermail 2.1.8 : Tue May 17 2005 - 00:02:57 EDT